Solve each formula for the indicated letter. Assume that all variables represent nonnegative numbers.
A = pi*r^2 + pi*rs, for r (Surface area of a cone)

Answer by stanbon(75887)

Solve each formula for the indicated letter. Assume that all variables represent nonnegative numbers.
A = pi*r^2 + pi*rs, for r (Surface area of a cone)
---
Rearrange:
pi*r^2 + pi*rs - A = 0
---
Divide thru by "pi":
r^2 + sr - A/pi = 0
---
Quadratic Formula:
r = [-s +- sqrt[s^2 - 4*1*(-A/pi)]/2
----
r = [-s +- sqrt[s^2 +[4A/pi]]]/2
---------------------------------------
Cheers,
Stan H.
=========